Group 2 Scenario:

You are working in a community outpatient clinic where you perform the intake assessment on R.M., a 38-year-old woman who is attending graduate school and is very sedentary. Her chief complaint is overwhelming fatigue that is not relieved by rest. She is so exhausted that she has difficulty walking to classes and trouble concentrating when studying. She reports a recent weight gain of 15 pounds over 2 months without clear changes in her dietary habits. Her face looks puffy, she has experienced excessive hair loss, and her skin is dry and pale. She complains of generalized body aches and pains with frequent muscle cramps and constipation. You notice she is dressed inappropriately warm for the weather.
Vital Signs
Blood pressure 142/84 mm Hg
Heart rate 52 beats/min
Respiratory rate 12 breaths/min
Temperature 96.8 ° F (36 ° C)

Laboratory Test Results
TSH 20.9 mU/L (2-10 mU/L)
TRH 18.8 ng/dL (2-10 ng/dL)
T 3 24 mU/L (70-205 ng/dL)
Free T 4 0.2 ng/dL (0.8-2.4 ng/dL)

1. What are the top priority assessments (specific to this patient)? Include at least 3 with rationale.


2. Interpret R.M.'s laboratory results.


3. The family practitioner prescribes levothyroxine (Synthroid) 1.7 mcg/kg body weight/day. Patient weighs 130 pounds. What is the daily dose in milligrams?


4. What teaching will you review regarding medication? List 5 key teaching points with details.


5. Why are we concerned with Myxedema coma? What teaching will you provide for prevention of myxedema coma?
Answer #1

1. The top priority assessments specific to the patient are :-

- excessive weight gain ( weight gain will lead to breathing difficulty and affect her performance )

- excessive fatigue

- generalized pain

2. TSH 20.9 mU/L (2-10 mU/L)

TRH 18.8 ng/dL (2-10 ng/dL)

T 3 24 mU/L (70-205 ng/dL)

Free T 4 0.2 ng/dL (0.8-2.4 ng/dL)

As per the laboratory results , she has Hypothyroidism due to which she has these sign and symptoms .

3. Ordered dose of levothyroxine is 1.7mcg / kg

weight of the patient is 130 Pounds

1.7mcg = 0.0017 mg

130pounds = 130× 0.45 = 58.5 kg

The daily dose in milligrams = 0.0017× 58.5 = 0.09945 mg

4. Nursing implications

-assess apical pulse & BP prior to & periodically during therapy

-assess for tachyarrhythmias & chest pain

-monitor thyroid function studies

-watch for OD manifested as hyperthyroidism

# Client teaching

-take med as directed

-instruct client on how to assess their own pulse

-hold dose if pulse >100 bpm

-explain therapy is life long

-advise patient to take tablets with water

-take on an empty stomach